# Exact Exploratory Bi-factor Analysis: A Constraint-based Optimisation Approach

### v1

@@ C.1 Proof of Theorem 1
where $\bm{\lambda}_{1}=\Lambda[\mathcal{B}^{*}_{g_{1}},\{1\}]$, $\bm{\lambda}_{g_{1}}=\Lambda[\mathcal{B}^{*}_{g_{1}},\{1+g_{1}\}]$, $\bm{\lambda}^{*}_{1}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1\}]$, $\bm{\lambda}^{*}_{g_{1}}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1+g_{1}\}]$, $\bm{\lambda}_{2,i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1\}]$, $\bm{\lambda}^{*}_{2,i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1\}]$, $\bm{\lambda}^{*}_{g_{2},i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1+g_{2}\}]$ for $i=1,2$, $\bm{\lambda}_{2,g_{1}}=\Lambda[\mathcal{B}^{*}_{g_{2},1},\{1+g_{1}\}]$, $\bm{\lambda}_{2,g^{\prime}_{2}}=\Lambda[\mathcal{B}^{*}_{g_{2},2},\{1+g^{\prime}_{2}\}]$, $\rho_{1,2}=\phi_{1+g^{\prime}_{1},1+g^{\prime}_{2}}$ and $\rho^{*}_{1,2}=\phi^{*}_{1+g^{\prime}_{1},1+g_{2}}$.
For the second case, there exists some $\alpha$ such that $\bm{\lambda}_{1}=\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}}$ and $\bm{\lambda}_{g_{1}}=\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}}$. Since the $\mathcal{B}_{g_{2}}^{*}$ rows and the $\mathcal{B}_{g_{2}}^{*}$ columns of $\Lambda\Phi(\Lambda)^{\top}$ and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$ have rank 2, under the bi-factor structure of the second case, we have $(\bm{\lambda}_{2,1},\bm{\lambda}_{2,g_{1}},\bm{\lambda}^{*}_{2,1},\bm{\lambda}^{*}_{g_{2},1})$ has rank 1 and $(\bm{\lambda}_{2,2},\bm{\lambda}_{2,g^{\prime}_{2}},\bm{\lambda}^{*}_{2,2},\bm{\lambda}^{*}_{g_{2},2})$ has rank 1 and $\bm{\lambda}^{*}_{g_{2},2}\neq\mathbf{0}$. We then assume that $\lambda^{*}_{2,2}=k_{1}\bm{\lambda}^{*}_{g_{2},2}$, $\bm{\lambda}_{2,2}=k_{2}\bm{\lambda}^{*}_{g_{2},2}$ and $\bm{\lambda}_{2,g^{\prime}_{2}}=k_{3}\bm{\lambda}^{*}_{g_{2},2}$. For the $\mathcal{B}^{*}_{g_{2},2}$ rows and the $\mathcal{B}^{*}_{g_{2},2}$ columns, we have $1+k_{1}^{2}=k_{2}^{2}+k_{3}^{2}$. For the $\mathcal{B}^{*}_{g_{1}}$ rows and the $\mathcal{B}^{*}_{g_{2},2}$ columns, we have
$\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{2,2})^{\top}+\rho^{*}_{1,2}\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{2},2})^{\top}$ (C.5)
$=$ $\bm{\lambda}_{1}(\bm{\lambda}_{2,2})^{\top}+\rho_{1,2}\bm{\lambda}_{g_{1}}(\bm{\lambda}_{2,g^{\prime}_{2}})^{\top}$
$=$ $k_{2}(\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}})(\bm{\lambda}^{*}_{g_{2},2})^{\top}+k_{3}(\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}})(\bm{\lambda}^{*}_{g_{2},2})^{\top}$
$=$ $(k_{2}\cos{\alpha}+k_{3}\sin{\alpha})\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{g_{2},2})^{\top}+(k_{3}\cos{\alpha}-k_{2}\sin{\alpha})\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{2},2})^{\top}.$
@@
For the first case, there exists some $\alpha,\beta$ such that $\bm{\lambda}_{1}=\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{g_{1}}=\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{2}=\cos{\beta}\bm{\lambda}^{*}_{2}-\sin{\beta}\bm{\lambda}^{*}_{g_{2}}$ and $\bm{\lambda}_{g_{2}}=\sin{\beta}\bm{\lambda}^{*}_{2}+\cos{\beta}\bm{\lambda}^{*}_{g_{2}}$. We then have the following equation
$\begin{pmatrix}\cos{\alpha}&\sin{\alpha}\\ -\sin{\alpha}&\cos{\alpha}\end{pmatrix}\begin{pmatrix}1&0\\ 0&\rho_{12}\end{pmatrix}\begin{pmatrix}\cos{\beta}&-\sin{\beta}\\ \sin{\beta}&\cos{\beta}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&\rho^{*}_{12}\end{pmatrix},$ (C.6)
which leads to $1=\cos{\alpha}\cos{\beta}+\rho^{*}_{12}\sin{\alpha}\sin{\beta}$. Since $|\rho^{*}_{12}|<1$, we have $\cos{\alpha}\cos{\beta}=1$ and $\sin{\alpha}\sin{\beta}=0$. Without loss of generation, we assume $\cos{\alpha}=\cos{\beta}=1$. Then we have $\bm{\lambda}_{1}=\bm{\lambda}^{*}_{1}$, $\bm{\lambda}_{g_{1}}=\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{2}=\bm{\lambda}^{*}_{2}$, $\bm{\lambda}_{g_{2}}=\bm{\lambda}^{*}_{g_{2}}$ and $\rho_{12}=\rho_{12}^{*}$.
@@
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}_{3}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{3}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,3}\\ 0&\rho_{1,3}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{3}^{\top}\\ \bm{\lambda}_{g_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{3}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}^{*}_{3}&\mathbf{0}&\bm{\lambda}^{*}_{g_{3}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,3}^{*}\\ 0&\rho_{1,3}^{*}&1\end{pmatrix}\begin{pmatrix}(\bm{\lambda}_{1}^{*})^{\top}&(\bm{\lambda}_{3}^{*})^{\top}\\ (\bm{\lambda}_{g_{1}}^{*})^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&(\bm{\lambda}_{g_{3}}^{*})^{\top},\end{pmatrix}$ (C.7)
@@
$\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{3})^{\top}+\rho^{*}_{1,3}\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{3}})^{\top}=\bm{\lambda}_{1}(\bm{\lambda}_{3})^{\top}+\rho_{1,3}\bm{\lambda}_{g_{1}}(\bm{\lambda}_{g^{\prime}_{3}})^{\top}$ (C.8)
@@
Since we have proved $\bm{\lambda}^{*}_{1}=\bm{\lambda}_{1}$ and $\bm{\lambda}^{*}_{g_{1}}=\bm{\lambda}_{g_{1}}$, we then have $\bm{\lambda}^{*}_{3}=\bm{\lambda}_{3}$, $\bm{\lambda}^{*}_{g_{3}}=\bm{\lambda}_{g^{\prime}_{3}}$ or $\bm{\lambda}^{*}_{g_{3}}=-\bm{\lambda}_{g^{\prime}_{3}}$ and $\rho^{*}_{1,3}=\rho_{1,3}$ or $\rho^{*}_{1,3}=-\rho_{1,3}$. Then, the proof is completed.
C.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
For any extended bi-factor model, we denote by $\mathcal{Q}_{g}=\{i\in\mathcal{B}_{g}|\lambda_{i,g+1}\neq 0\}$ and $H=\{g|\Lambda[\mathcal{B}_{g},\{1,1+g\}]\mbox{~{}has column rank 2}\}$. The following condition is the necessary condition for the identifiability of the extended bi-factor model under a known bi-factor structure, as proposed in Theorem 3 of Fang et al., (2021).
@@ C.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
$|H|\geq 2$; $|\mathcal{Q}_{g}|\geq 2$ for $1,\ldots,G$ and $|\mathcal{Q}_{g}|\geq 3$ for $g$ that satisfies $\phi_{1+g,1+g^{\prime}}=0$ for all $g^{\prime}\neq g$.
@@
$|H^{*}|\geq 3$. In addition, there exist $g1,g2,g3\in H^{*}$ such that (1) $|Q_{g1}^{*}|\geq 3$, $|Q_{g2}^{*}|\geq 3$ and $|Q_{g3}^{*}|\geq 3$, (2) any 2 rows of $\Lambda^{*}[\mathcal{B}_{g1}^{*},\{1+g_{1}\}]$ are linearly independent.
@@
Condition 5 is easy to check in practice and $\widehat{\Lambda}$ in Section 4 satisfies Condition 5.
@@
Suppose the true bi-factor structure follows $\widehat{\Lambda}$ in Section 4. Let $\Lambda^{*}$, $\Phi^{*}$ and $\Psi^{*}$ be the true parameters such that Condition 4and 5 are satisfied. For any parameters $\Lambda$, $\Phi$ and $\Psi$ that satisfy Condition 4 and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}=\Lambda\Phi(\Lambda)^{\top}+\Psi$, there exists a diagonal sign-flip matrix $D\in\mathcal{D}$ and a permutation matrix $P\in\mathcal{P}$ such that $\Lambda=\Lambda^{*}PD$, $\Phi=DP^{\top}\Phi^{*}PD$ and $\Psi^{*}=\Psi$.

### v2

@@ G.1 Proof of Theorem 1
where $\bm{\lambda}_{1}=\Lambda[\mathcal{B}^{*}_{g_{1}},\{1\}]$, $\bm{\lambda}_{g_{1}}=\Lambda[\mathcal{B}^{*}_{g_{1}},\{1+g_{1}\}]$, $\bm{\lambda}^{*}_{1}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1\}]$, $\bm{\lambda}^{*}_{g_{1}}=\Lambda^{*}[\mathcal{B}^{*}_{g_{1}},\{1+g_{1}\}]$, $\bm{\lambda}_{2,i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1\}]$, $\bm{\lambda}^{*}_{2,i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1\}]$, $\bm{\lambda}^{*}_{g_{2},i}=\Lambda[\mathcal{B}^{*}_{g_{2},i},\{1+g_{2}\}]$ for $i=1,2$, $\bm{\lambda}_{2,g_{1}}=\Lambda[\mathcal{B}^{*}_{g_{2},1},\{1+g_{1}\}]$, $\bm{\lambda}_{2,g^{\prime}_{2}}=\Lambda[\mathcal{B}^{*}_{g_{2},2},\{1+g^{\prime}_{2}\}]$, $\rho_{1,2}=\phi_{1+g_{1},1+g^{\prime}_{2}}$ and $\rho^{*}_{1,2}=\phi^{*}_{1+g_{1},1+g_{2}}$.
For the second case in (G.4), there exists some $\alpha$ such that $\bm{\lambda}_{1}=\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}}$ and $\bm{\lambda}_{g_{1}}=\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}}$. Since the $\mathcal{B}_{g_{2}}^{*}$ rows and the $\mathcal{B}_{g_{2}}^{*}$ columns of $\Lambda\Phi(\Lambda)^{\top}$ and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}$ have rank 2, under the bi-factor structure of the second case, we have that $(\bm{\lambda}_{2,1},\bm{\lambda}_{2,g_{1}},\bm{\lambda}^{*}_{2,1},\bm{\lambda}^{*}_{g_{2},1})$ has rank 1 and $(\bm{\lambda}_{2,2},\bm{\lambda}_{2,g^{\prime}_{2}},\bm{\lambda}^{*}_{2,2},\bm{\lambda}^{*}_{g_{2},2})$ has rank 1. Noticing that $\bm{\lambda}^{*}_{g_{2},1}\neq\mathbf{0}$, we assume that $\lambda^{*}_{2,1}=k_{1}\bm{\lambda}^{*}_{g_{2},1}$, $\bm{\lambda}_{2,1}=k_{2}\bm{\lambda}^{*}_{g_{2},1}$ and $\bm{\lambda}_{2,g_{1}}=k_{3}\bm{\lambda}^{*}_{g_{2},1}$. For the $\mathcal{B}^{*}_{g_{2},1}$ rows and the $\mathcal{B}^{*}_{g_{2},1}$ columns of (G.4), we have $1+k_{1}^{2}=k_{2}^{2}+k_{3}^{2}$. For the $\mathcal{B}^{*}_{g_{1}}$ rows and the $\mathcal{B}^{*}_{g_{2},1}$ columns of (G.4), we have
$\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{2,1})^{\top}+\rho^{*}_{1,2}\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{2},1})^{\top}$ (G.5)
$=$ $\bm{\lambda}_{1}(\bm{\lambda}_{2,1})^{\top}+\bm{\lambda}_{g_{1}}(\bm{\lambda}_{2,g_{1}})^{\top}$
$=$ $k_{2}(\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}})(\bm{\lambda}^{*}_{g_{2},1})^{\top}+k_{3}(\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}})(\bm{\lambda}^{*}_{g_{2},1})^{\top}$
$=$ $(k_{2}\cos{\alpha}+k_{3}\sin{\alpha})\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{g_{2},1})^{\top}+(k_{3}\cos{\alpha}-k_{2}\sin{\alpha})\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{2},1})^{\top}.$
@@
For the first case in (G.3), there exists some $\alpha,\beta$ such that $\bm{\lambda}_{1}=\cos{\alpha}\bm{\lambda}^{*}_{1}-\sin{\alpha}\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{g_{1}}=\sin{\alpha}\bm{\lambda}^{*}_{1}+\cos{\alpha}\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{2}=\cos{\beta}\bm{\lambda}^{*}_{2}-\sin{\beta}\bm{\lambda}^{*}_{g_{2}}$ and $\bm{\lambda}_{g_{2}}=\sin{\beta}\bm{\lambda}^{*}_{2}+\cos{\beta}\bm{\lambda}^{*}_{g_{2}}$. We then have the following equation
$\begin{pmatrix}\cos{\alpha}&\sin{\alpha}\\ -\sin{\alpha}&\cos{\alpha}\end{pmatrix}\begin{pmatrix}1&0\\ 0&\rho_{12}\end{pmatrix}\begin{pmatrix}\cos{\beta}&-\sin{\beta}\\ \sin{\beta}&\cos{\beta}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&\rho^{*}_{12}\end{pmatrix},$ (G.6)
which leads to $1=\cos{\alpha}\cos{\beta}+\rho_{12}\sin{\alpha}\sin{\beta}$. Since $|\rho_{12}|<1$, we have $\cos{\alpha}\cos{\beta}=1$ and $\sin{\alpha}\sin{\beta}=0$. Without loss of generation, we assume $\cos{\alpha}=\cos{\beta}=1$. Then we have $\bm{\lambda}_{1}=\bm{\lambda}^{*}_{1}$, $\bm{\lambda}_{g_{1}}=\bm{\lambda}^{*}_{g_{1}}$, $\bm{\lambda}_{2}=\bm{\lambda}^{*}_{2}$, $\bm{\lambda}_{g_{2}}=\bm{\lambda}^{*}_{g_{2}}$ and $\rho_{12}=\rho_{12}^{*}$.
@@
$\begin{pmatrix}\bm{\lambda}_{1}&\bm{\lambda}_{g^{\prime}_{1}}&\mathbf{0}\\ \bm{\lambda}_{3}&\mathbf{0}&\bm{\lambda}_{g^{\prime}_{3}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,3}\\ 0&\rho_{1,3}&1\end{pmatrix}\begin{pmatrix}\bm{\lambda}_{1}^{\top}&\bm{\lambda}_{3}^{\top}\\ \bm{\lambda}_{g^{\prime}_{1}}^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&\bm{\lambda}_{g^{\prime}_{3}}^{\top}\end{pmatrix}=\begin{pmatrix}\bm{\lambda}^{*}_{1}&\bm{\lambda}^{*}_{g_{1}}&\mathbf{0}\\ \bm{\lambda}^{*}_{3}&\mathbf{0}&\bm{\lambda}^{*}_{g_{3}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&\rho_{1,3}^{*}\\ 0&\rho_{1,3}^{*}&1\end{pmatrix}\begin{pmatrix}(\bm{\lambda}_{1}^{*})^{\top}&(\bm{\lambda}_{3}^{*})^{\top}\\ (\bm{\lambda}_{g_{1}}^{*})^{\top}&\mathbf{0}^{\top}\\ \mathbf{0}^{\top}&(\bm{\lambda}_{g_{3}}^{*})^{\top},\end{pmatrix}$ (G.7)
@@
$\bm{\lambda}^{*}_{1}(\bm{\lambda}^{*}_{3})^{\top}+\rho^{*}_{1,3}\bm{\lambda}^{*}_{g_{1}}(\bm{\lambda}^{*}_{g_{3}})^{\top}=\bm{\lambda}_{1}(\bm{\lambda}_{3})^{\top}+\rho_{1,3}\bm{\lambda}_{g_{1}}(\bm{\lambda}_{g^{\prime}_{3}})^{\top}$ (G.8)
@@
Since we have proved $\bm{\lambda}^{*}_{1}=\bm{\lambda}_{1}$ and $\bm{\lambda}^{*}_{g_{1}}=\bm{\lambda}_{g_{1}}$, we then have $\bm{\lambda}^{*}_{3}=\bm{\lambda}_{3}$, $\bm{\lambda}^{*}_{g_{3}}=\bm{\lambda}_{g^{\prime}_{3}}$, $\rho^{*}_{1,3}=\rho_{1,3}$ or $\bm{\lambda}^{*}_{g_{3}}=-\bm{\lambda}_{g^{\prime}_{3}}$, $\rho^{*}_{1,3}=-\rho_{1,3}$.
Now, since $\Lambda$ has $G$ group factors, according to the previous proof, each variable belonging to $\mathcal{B}_{i}^{*}$ loads on a unique group factor according to $\Lambda$ and the loadings of the general factor and the group factors are determined up to a sign flip for $i=1,\ldots,G$. Thus, there exist a diagonal sign-flip matrix $D\in\mathcal{D}$ and a permutation matrix $P\in\mathcal{P}$ such that $\Lambda=\Lambda^{*}PD$. It is straightforward to further check that $\Phi=DP^{\top}\Phi^{*}PD$. Thus, the proof is completed.
G.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
For any matrix $A$, we use $\text{rank}(A)$ to denote the rank of $A$. The following condition is a necessary condition for the identifiability of the extended bi-factor model under a known bi-factor structure, as proposed in Theorem 3 of Fang et al., (2021).
@@ G.2 Identifiability of Estimated Bi-factor Structure in Real Data Example
$|\mathcal{B}_{g}|\geq 2$ for all $g=1,\ldots,G$.
@@
For any $m\times n$ dimensional submatrix of $\Phi[\{2,\ldots,1+G\},\{2,\ldots,1+G\}]$, $1\leq m,n\leq G$, it’s rank is $\min(m,n)$.
Condition 6.
For any g such that $|\mathcal{B}^{*}_{g}|\geq 3$, any 2 rows of $\Lambda^{*}[\mathcal{B}_{g}^{*},\{1,1+g\}]$ are linearly independent.
@@
Condition 5 restricts that the correlation matrix of group factors does not degenerate. In Theorem 2, we restrict the parametric space of $\Phi$ to the space satisfying Condition 5. We note that $\widehat{\Lambda}$ in Section 4 satisfies Condition 5. Condition 6 is easy to check in practice and $\widehat{\Lambda}$ in Section 4 satisfies Condition 6.
@@
Suppose the true bi-factor structure follows $\widehat{\Lambda}$ in Section 4. Let $\Lambda^{*}$, $\Phi^{*}$ and $\Psi^{*}$ be the true parameters such that Conditions 4 -6 are satisfied. For any parameters $\Lambda$, $\Phi$ and $\Psi$ that satisfy Conditions 4 and 5 and $\Lambda^{*}\Phi^{*}(\Lambda^{*})^{\top}+\Psi^{*}=\Lambda\Phi(\Lambda)^{\top}+\Psi$, there exists a diagonal sign-flip matrix $D\in\mathcal{D}$ and a permutation matrix $P\in\mathcal{P}$ such that $\Lambda=\Lambda^{*}PD$, $\Phi=DP^{\top}\Phi^{*}PD$ and $\Psi^{*}=\Psi$.